OpenStudy (anonymous):

Let f(x) = x^2 – 16. Find f^–1(x).

OpenStudy (anonymous):

Let f(x) = y So, \[y = x^2 - 16\] \[x^2 = y + 16\] \[x = \sqrt{y+16}\] Just swap x and y now: \[y = \sqrt{x + 16}\] This is \(f^{-1}(x)\)..

OpenStudy (anonymous):

@waterineyes...why not ! \[-\sqrt{y+16}\]

OpenStudy (anonymous):

Yes it can be but I here only consider the positive value of root..

OpenStudy (anonymous):

\[\pm 4\sqrt{}x\] \[\pm \sqrt{+ 16} \] \[x ^{2} \div 16\] \[1\div x^{2} - 16\]

OpenStudy (anonymous):

one of those have to be the answers

OpenStudy (anonymous):

then, they mean 1/f(x)...it's number 4 !

OpenStudy (anonymous):

the last question ?

OpenStudy (anonymous):

go on if you have more questions :)

OpenStudy (anonymous):

lol so the answer is the fourth . correct?

OpenStudy (anonymous):

yeaaah...\[1/(x^2-16)\]

OpenStudy (anonymous):

ohh ok thank youu (:

OpenStudy (anonymous):

yw ! anytime !